AGw Commissioning with IRTC 01-02 November 2008 MST

Observers: D.Thompson, J. Hill on mtn at beginning, J.Storm from Potsdam
Software Support: S.Hooper, N. Cushing C.Biddick on call as needed.
Telescope Operator: D. Gonzalez Huerta

Summary

Large (108 arcsec) pointing offset from previous night (back to starting point from several nights ago). Source unknown. Resetting the encoders did not change the result.

Checked IRTC orientation (it differs from LUCIFER by ~1.2 degrees) and rotator center (it is offset by ~8-9 pixels from last time it was mounted). Some "failed" presets are from my iraf scripts not handling warnings correctly.

Rotator twice went into fault mode. An rcready did not immediately clear this, but a subsequent one did. No source of problem identified.

Worked on AGw centering of the hotspot on the rotator mechanical axis. The tests were giving unstable results. It appears the ~1 arcsec jumps affect this as well. Centering not completed, but we are within 1 arcsec or so.

Details (times reported in UT below)

startup IRS- IRTC+ AGW- (note, no longer any need to start IRS or AGW from iraf script)

01:40 pointto BS9183 gs=0 (AGw on axis) failed - MCSPU error returned - ERROR : no polys are being received!
  • restarting PCS

01:47 pointto BS9183 gs=0 PAM- Tel="ACQUIRE" - success, correct pointing off ~20", but use a brighter star...

pointto WT10_219 gs=0 PAM- Tel="ACQUIRE" - dropped out of acquire mode 02:04, not obvious why, but we did adjust the exposure time on the acquisition images just before it failed. GCS got a "stop guiding" command at 19:03:17. Happened again at 19:10:32, on modification of CA at EL=59. John thinks this incurs a large enough AZ offset that the telescope slews and IIF sees this and drops out of the ACQUIRE mode. See IssueTrak#1796

Resend preset...

Used spiral search to find star. It was ~120 arcsec off the previous night's pointing! IE=65. CA=-24 from Nov01 (end of last night), tonight it is IE=-43 CA=-31. Active optics were cleared at the start of tonight, and we are on the strip encoders. How are the zero points set for the strip encoders?

Confirmed pointing OK on another nearby pointing star. Lost ~1 hour correcting pointing due to the large change from last night (108 arcsec in IE and only 7 arcsec in CA).

02:45 pointto BS9183 gs=0 PAM- Tel="ACTIVE", successful. GCS is guiding, WFSing on axis. Allow to collimate.

Updated IRTC headers to include telescope temperature sensor data. Confirmed.

02:58 Telescope collimated. WFS sending ~120nm corrections per cycle. Guider reports 0.8" FWHM.

03:01 pointto BS9183 gs=32 Tel="ACTIVE" - preset failed around 03:03, rotator in wait_OPR mode. No failure message sent back through IIF, from the IRTC side the preset appears to be hung. Calling D.Ashby @ 8:07MST (14 minutes). No obvious solution?

03:21 Resetting the encoders. DAshby thought it a good idea to do this each night and that the 108 arcsec jump was due to not resetting the encoders. Encoders needed 3 tries to reset. Done at 03:55.

03:56 pointto WT10_204 gs=0 PAM- Tel="ACQUIRE" - succeeded. IE=-43, CA=-31 same as set earlier. Star in center of acquisition field.

04:00 pointto BS9186 gs=9 Tel="ACTIVE" PAM+ - succeeded! Somewhat out of collimation. First WFS=1093nm, second=862. Aberrations mostly symmetric (focus, spherical?).

Check IRTC orientation the quick way... preset in TRACK mode, PA=0
image coords of star offset
041933 140.26 160.87 0.0 0.0
042044 207.41 158.65 7.0 0.0
042213 66.79 161.51 -7.0 0.0
042343 132.71 110.10 0.0 5.0
042453 134.62 207.60 0.0 -5.0
042558 132.76 160.65 0.0 0.0

Change IRTCBFG LEFTZEROPOINT to -186.29 in config file, reauthorize IRTC on IIF GUI.

04:43 Rotator in fault mode again. Calling D.Ashby 9:45MST (23 min). First "rcready lfbg" did not clear the fault state, but a second one a couple minutes later did clear the fault state.

Re-check IRTC orientation after update...
image coords of star offset
051940 188.09 148.98 0.0 0.0
052005 257.96 147.80 7.0 0.0
052027 120.44 149.71 -7.0 0.0

In ACTIVE mode
052617 152.95 150.40 0.0 0.0
052725 40.81 116.91 12.0 0.0
052903 265.59 181.02 -12.0 0.0
053204 264.86 180.74 -12.0 0.0
053501 41.56 116.49 12.0 0.0
6.719 564.001

Not at PA=0, duh!

pointto BS9186 gs=6 pa=0 Tel="ACTIVE"
054807 -- 0.0 0.0
054944 42.12 158.01 -12.0 0.0
055104 275.40 156.44 12.0 0.0

We are below 1 degree on the alignment between IRTC and AGw (I get 0.38 degrees).

060208-060635 PA=0 and 180 data to recheck rotator center:

218.18 144.31 84.73 128.10 151.28 136.21
84.73 128.10 217.83 144.61 151.28 136.36
219.97 144.37 84.61 128.18 152.29 136.28

Rotator center on IRTC: 151.62 136.28 This is about 8 pixels offset from previous installation

Check DDViewer for instrument PA in IRTC headers...pcs.preset.pa is correct. This gives the PA of the Y axis of the IRTC projected on the sky.

Several presets fail due to "couldn't adjust exposure time to meet low counts requirement. star too dim or guide rate too high." Latest was started at UT06:52:08 or shortly thereafter. But guiding and WFS both started, so this is an incorrect failure sent back through IIF. It looks like this is an incorrect handling of warnings by my LBTtools/IRTC package, will update it tomorrow.

06:57 Seeing is getting worse. The guide cam was below 1 arcsec until now, but EL<40. Moving to BS9105. EL~63 and transiting now.

Working on center script, corrected the trig and it now seems to work at any PA. You need to take an image, run the center script on its timestamp.

AGw calibrations

Jesper joined us at 08:00

08:20 Sent preset to M34 astrometric field, GUIDE mode, PA=0, allow to collimate.

Images in /Repository/20081102_IRTC/guider*.fits

5s exposure before starting rotator: /Repository/20081102_IRTC/guider000002.fits

Check rotator mechanical center in (nominally) on-axis AGW probe: /Repository/20081102_IRTC/guider000003.fits
Center at 287,237 (DThompson), 288,236 (JStorm).

AIP_L.cfg has incorrect hotspot coordinates (292,255)! They should be HOTSPOT x=289, y=252 (determined on 14-15 Oct). This needs to be corrected in SVN to prevent future overwrites. Jesper also updated the oacontrol.conf file on the AGw computer.

AGw: stopAGW, oacserver stopped and restarted, and startAGW. Multiple selectAGW's failed to find the unit. Restarting GCS at 09:27, that worked.

guider000005 = 5s
probe moved slightly to the newly redefined on-axis position
guider000006 = 60s while rotating
Center at 282 230 (DThompson), 284 230 (JStorm).
Files updated, software stopped and restarted as before.

Got another "couldn't adjust exposure time to meet low counts requirement" message with a failed preset (the result code back from the IRS is 2, is this just a warning? I'm reporting failure for anything other than a zero in the iraf scripts), UT10:00. See guiderimage011776, TIME-OBS 09:59:05.43, 4.6 seconds exposure and the star is ~380 counts above background. GCS acquisition logic needs tuning.

guider000007 = 5s, TRACK mode
probe moved slightly to the redefined on-axis position
guider000008 = 60s while rotating - Center at 284 244 (DThompson), 284 245 (JStorm)
guider000010 = 60s while rotating - Center at 285 268 (DThompson), 284 265 (JStorm)
guider000012 = 60s while rotating - Center at 283 253 (DThompson), 284 250 (JStorm)
guider000014 = 60s while rotating - Center at 279 264 (DThompson) - offsets and results not making sense.

Will try rotating back and forth 2x each direction with no other changes to see if the rotator is doing this.
guider000015 rotator -70 -> 200 degrees, center at 278 252
guider000016 rotator 200 -> -70 degrees, center at 275 252
guider000017 rotator -70 -> 200 degrees, center at 279 251
guider000018 rotator 200 -> -70 degrees, center at 276 252

recollimate, then retest

guider000019 rotator -68 -> 202 degrees, center at 279 252
guider000020 rotator 202 -> -68 degrees, center at 276 251
guider000021 rotator -68 -> 202 degrees, center at 282 253
guider000022 rotator 202 -> -68 degrees, center at 276 252

It appears we are affected by the jumps here as well.
recollimate, then retest (oacserver config updated again, stuff stopped and restarted)

guider000023 rotator -63 -> 207 degrees, center at 290 261
guider000024 rotator 207 -> -63 degrees, center at 287 263
guider000025 rotator -63 -> 207 degrees, center at 290 263
guider000026 rotator 207 -> -63 degrees, center at 286 264

one more correction attempt:

guider000028 rotator -60 -> 210 degrees, center at 282 251
guider000029 rotator 210 -> -60 degrees, center at 289 251
guider000030 rotator -60 -> 210 degrees, no star visible, sky is too bright

UT12:50 closing - end of night.

-- DavidThompson - 02 Nov 2008
Topic revision: r5 - 03 Nov 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback